The equation for the formation of hydrogen iodide
Answer #1

H2 + I2 ---> 2HI I 0.1800 0.1800 C:-X E: (01800-х) (0.1800-х) +2x Кр = P(H1)2/P(H2)P(12) 69.0 = (2x)2 (0.1 800-х)2 69.0 =